OpenStudy (anonymous):

dx/dt=2y-x dy/dt=e^x+y solve y=f(x)

OpenStudy (anonymous):

can we just use the fact that it is exact?

OpenStudy (jamesj):

Hm ... are you sure the second equation is e^x not e^t?

OpenStudy (mr.math):

Let me start by substituting \(2y=x'+x \implies 2y'=x''+1\). We can rewrite the second DE as \(x''+1=2e^x+x'+x \implies x''-x'-x=2e^x-1\).

OpenStudy (jamesj):

@imran, no you most definitely cannot do that because y is a function of x and vice versa so your integrations are incorrect. @Mr.M: you've got it.

OpenStudy (mr.math):

But e^x is a problem.

OpenStudy (anonymous):

This is phase plane problem , btw

OpenStudy (mr.math):

I don't even know what phase plane problems are!

OpenStudy (anonymous):

b/c we can divide them dx/dt=2y-x dy/dt=e^x+y dy/dx= e^x+y ----- 2y-x (2y-x)dy= (e^x+y) dx (e^x+y) dx +(-2y+x)dy= 0 exact e^x+yx -y^x=C